# Trade-off of Security, Latency, and Throughput of the Nakamoto Consensus

### v3

@@ 2.3. Adversary and Propagation Delay
A block is either an A-block or an H-block. The parent of an H-block must be credible when the H-block is mined.
An A-block (resp. H-block) is a mathematical concept that represents a block mined by an adversarial (resp. honest) miner. According to Definitions 5 and 9, an H-block must be higher than all public blocks at the time of its mining. There is no such height constraint on an A-block. In fact, this is the only constraint on block heights. In Fig. 1, we distinguish each A-block using a pair of horns on top. Blocks 2 and 7 are A-blocks, while all the other blocks are H-blocks.
Definition 10 (transaction’s inclusion).
Suppose a transaction becomes public at time $s$. If block $b$ is the first H-block mined after $s$, then chain $b$ must include the transaction.
The propagation delays of transactions, blocks, and chains depend on the network conditions and may also be manipulated by the adversary. To establish a strong safety guarantee, we adopt the widely used $\Delta$-synchrony model.
@@
Definition 10 guarantees that once a transaction becomes public, it will be promptly included in at least one chain. Together, Definitions 10 and 11 imply that, if a transaction becomes public at $s$, no H-block mined before time $s-\Delta$ includes it, while it may be included in an A-block mined before $s-\Delta$.
@@
Let $A_{s,t}$ denote the number of A-blocks mined during $(s,t]$. Let $H_{s,t}$ denote the number of H-blocks mined during $(s,t]$. Evidently, $(A_{0,t}+H_{0,t})_{t\geq 0}$ is a Poisson point process with rate $a+h$. As a convention, let $A_{s,t}=H_{s,t}=0$ for all $s>t$. Since the probability that any two blocks are mined at exactly the same time is equal to zero, we assume without loss of generality that all blocks are mined at different times. Hence $t_{0}<t_{1}<t_{2}<\dots$.
3. Model Sufficiency
3.1. Sufficiency
@@
The prevailing model in this line of work (see, e.g., (garay2015bitcoin, ; pass2017analysis, ; pass2017rethinking, ; garay2017bitcoin, ; dembo2020everything, ; gazi2020tight, ; saad2021revisiting, ; neu2022longest, ; gazi2023practical, ; kiffer2023security, )) is based on the universally composable security framework (canetti2000universally, ). A complete description of the model, henceforth referred to as the view-based model, requires the definition of a protocol as a set of algorithms, the execution of those algorithms that is directed by an environment, which activates $n$ either honest or corrupted parties with individual views (pass2017analysis, , Sec. 2). The protocol’s execution is no simpler than a system of $n$ interactive Turing machines. Due to its complexity, the view-based model always assumes discrete time steps in execution.
@@
The model proposed in Sec. 2 of this paper is entirely self-contained and much simpler than the view-based model. In contrast to previous works as well as (li2021close, ; guo2022bitcoin, ), our model intentionally exclude miners and their views in definitions. In particular, with well-defined block mining processes, we are not concerned with the number of miners; neither do we keep track of which blocks an individual miner has mined or seen. Instead, our model keeps track of the set of public chains at every point in time, which represents the weakest possible view of any miner. In particular, every honest view is a superset of this weakest view, and conversely, the adversary in the view-based model can ensure at least one honest miner holds that weakest view.
@@
Relative to the view-based model under $\Delta$-synchrony, our model has no loss of generality in the sense that, for a given realization (or sample path) of the A- and H-block mining processes, if the view-based model can produce a certain set of chains, so can our model. This is because both models respect the same constraints for chain creation described in Definitions 9 and 11. For the latency-security analysis, our model is as powerful as the view-based model (if simpler). To elaborate, consider the event in the view-based model where, first, a chain in an honest view confirms a block on height $h$, and then another chain in an honest view confirms a different block on the same height, creating a safety violation. Our model can produce the same pair of chains, each of which is credible to the weakest view at the time it confirms a conflicting block. Conversely, if our model produces a pair of conflicting credible chains, then those chains can be used to convince two honest parties in the view-based model to confirm different blocks on the same height. Once we realize the role of the weakest view, there is no need to define individual views and their complicated interactions.
3.2. Operations
@@
For concreteness and practicality, we relate the definitions in Sec. 2 to the operations of miners (or nodes), who assemble transactions into blocks and chains by expending proof of work. The same miners also propagate those chains through a peer-to-peer network. The honest miners follow the longest-chain-fork-choice rule whereas the adversary can use an arbitrary (possibly randomized) strategy. At time $t$, each node has its own view, which includes all transactions, blocks, and chains the node has seen by $t$. A node’s view can only increase with time. The operations of each node is based on the entire history of this node’s view.
In general, an honest node will only provide goods and services for a transaction after it is considered committed. If an honest node first encounters a transaction at time $s$, it can deduce that the transaction will become public during the time window of $[s,s+\Delta]$. Here we define two commit rules corresponding to the two types of confirmation defined in Sec. 2. Specifically, if the transaction is included in a highest chain in this node’s view at time $s+r$ for some $r\geq t$, and the chain remains the highest at $s+r+\Delta$, then the node can infer that 1) the chain is credible and public at $s+r+\Delta$, and 2) the transaction has been confirmed by time $t$ in the mathematical model. In this case, the node commits the transaction by time $t$. Similarly, if block $d$ is the highest block in this node’s view at time $s+\Delta$, and there is a chain $c$ in the node’s view that remains the highest for a period of $\Delta$ or longer, which contains a block $b$ including the transaction, and $h_{c}\geq h_{b}+k-1$ as well as $h_{c}\geq h_{d}+k$, then the node can infer that the transaction has been confirmed by depth $k$ in the mathematical model. In this case, the node commits the transaction by depth $k$.
We will show that, regardless of the adversary’s strategy, violating the safety of a transaction implies that an “error event” concerning the two mining processes occurs. We then bound the probability of safety violation defined in Sec. 2, which implies a probabilistic guarantee for an honest node’s operational safety.
3.3. Delay Variations
@@
The seemingly rigid $\Delta$-synchrony model is quite practical. First, “light” nodes do not mine blocks or fork chains, so the delays they perceive are inconsequential to ledger safety as long as honest miners satisfy the delay requirement. Second, if some blocks of honest miners suffer longer delays than $\Delta$, we simply regard those blocks as adversarial blocks to preserve $\Delta$-synchrony. In addition, potential attacks like Erebus (Erebus, ) and SyncAttack (SyncAttack, ) seek to partition the peer-to-peer network. As long as those delay violations are independent of the block mining times, the honest and adversarial mining processes remain to be Poisson. If block propagation delays have a long-tailed distribution (decker2013information, ; fechner2022calibrating, ), we can use a delay bound that applies to a high percentage of blocks. The main results here continue to apply with all out-of-sync mining power treated as being adversarial.
In Definition 10, we have assumed that every public transaction gets included in a chain very soon. In a more general model, a transaction’s inclusion can be delayed due to censorship or a lack of space in new blocks. If this is the case, we modify the confirmation rules as follows: First, block $b$ which contains the target transaction becomes public. Then, if block $b$ is in a credible public chain $c$ with $h_{c}\geq h_{b}+k-1$, we say chain $c$ confirms the transaction by depth $k$. If block $b$ is in a credible public chain $c$ at time $t_{b}+r$ with $r\geq t$, we say chain $c$ confirms the transaction by time $t$. We note that the main theorems in the next section can be easily generalized to this case, although we omit the details in this paper.
4. Main Results
@@
Recall that $a$ and $h$ represent as the adversarial and honest mining rates, respectively, in the number of blocks mined per second. For confirmation by either time or depth, the following negative result is well understood (see, e.g., (pass2017analysis, ; dembo2020everything, ; gazi2020tight, )).
@@
(2) $\frac{1}{a}<\Delta+\frac{1}{h},$
then every transaction is unsafe in the sense that the private-mining attack on the transaction will violate its safety with probability 1.
@@
4.1. Bounds for Given Confirmation Depth
@@
(3) $b(\alpha,\eta)=\left(\frac{e^{4\gamma}\gamma(\eta-\alpha-\alpha\eta)}{\alpha\eta e^{\gamma}-(\alpha+\gamma)(\eta-\gamma)}\right)^{2}+1-\frac{\eta-\alpha-\alpha\eta}{\eta e^{\alpha}},$
@@
(4) $\begin{split}c&(\alpha,\eta)=\left(\ln\left(1+\frac{\gamma}{\alpha}\right)+\ln\left(1-\frac{\gamma}{\eta}\right)-\gamma\right)\\ &\cdot\left(1-\ln\left(1+\frac{\gamma}{\alpha}\right)\bigg{/}\ln\frac{\eta e^{\alpha}(e^{\eta}-1)}{\eta+\eta e^{\alpha}(e^{\eta}-1)-\alpha(1+\eta)}\right)^{-1}\end{split}$
@@
(5) $\gamma(\alpha,\eta)=\frac{2-\alpha+\eta-\sqrt{4+(\alpha+\eta)^{2}}}{2}.$
The following latency–security trade-off is the first set of non-trivial bounds that apply to all parameter values up to the ultimate fault tolerance.
Theorem 2 (Upper bounding the violation probability for given confirmation depth).
@@
(6) $\frac{1}{a}>\Delta+\frac{1}{h}.$
Given a sufficiently large $s$, the safety of a transaction which becomes public during $(s,s+2\Delta]$ and is confirmed by the depth of $k$ can be violated with a probability no higher than
(7) $b(a\Delta,h\Delta)e^{-c(a\Delta,h\Delta)\cdot k}$
@@
As long as (6) holds, every transaction is arbitrarily safe after sufficiently long confirmation latency (either in depth or time).
The pair of opposing inequalities (2) and (6) determine the exact boundary between the asymptotically achievable and unachievable latency–security trade-offs. The right hand side of (2) and (6) is exactly the expected inter-arrival time of the following type of blocks:
@@
By definition, pacers become public by their immediate successors and hence all pacers are on different heights. Since the expected pacer inter-arrival time is $\Delta+1/h$, pacers arrive at the rate of $h/(1+h\Delta)$ in the long run. As a consequence, the number of heights H-blocks occupy grows at least at that rate. This approach enables us to use the growth rate of pacers as a lower bound for the height increase of H-blocks.
@@
(8) $f(t)=\frac{1}{h}e^{-h(t-\Delta)}1_{\{t>\Delta\}}.$
We also consider a delayed renewal process in general, which is like an ordinary renewal process, except that the first arrival time (delay) may have a different distribution than the other inter-arrival times.
Lemma 4 ((Kroese_difference, , Theorem 3.4)).
@@
(10) $\mathcal{E}(r)=\frac{(1-r)(h-a-ha\Delta)}{h-e^{(1-r)a\Delta}(h+a-ar)r}.$
@@
(11) $e(i)$ $=\frac{1}{i!}\mathcal{E}^{(i)}(0),\qquad i=0,1,\dots$
@@
(13) $\delta$ $=\left(1-a\frac{1+h\Delta}{h}\right)^{+}e^{-a\Delta}.$
@@
(14) $F^{(k)}(n)=1-(1-\delta)\left(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\right)^{n+1}\bar{F}_{b}(n;k,(\epsilon+\delta-\epsilon\delta))$
@@
(15) $F(n)$ $=\lim_{k\to\infty}F^{(k)}(n)$
(16) $=1-(1-\delta)\left(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\right)^{n+1}.$
@@
Given a sufficiently large $s$, the safety of a transaction which becomes public during $(s,s+2\Delta]$ and is confirmed by the depth of $k$ can be violated with a probability no higher than
$\inf_{n}\bigg{\{}2-$ $\sum_{i=0}^{k-n}e(i)\sum_{j=0}^{k-n-i}e(j)\int_{0}^{\infty}F_{1}(k-n-i-j;a(t+6\Delta))$
(17) $\cdot f_{2}(t-(k)\Delta;k,h)\,dt-F^{(k)}(n-1)\bigg{\}}$
@@
Definition 14 (Private-mining attack on a target transaction).
The adversary delays the propagation of every H-block maximally (for $\Delta$ seconds). Starting from time 0, the adversary always mines in private to extend the highest chain which does not include the target transaction. The adversary makes public the private chain as soon as both of the following conditions hold: (1) the transaction has been confirmed by a chain of H-blocks; and (2) the private chain which does not contain the transaction is credible.
@@
Given a sufficiently large $s$, the private-mining attack violates the safety of a target transaction which becomes public during $(s,s+2\Delta]$ with at least the following probability:
@@
$\cdot\sum_{j=0}^{k-i}e(j)\bigg{(}f_{1}(0,h\Delta)\int_{0}^{\infty}F_{1}(k-(1+i+j);a(t-\Delta))$
$\cdot f_{2}(t-k\Delta;k,h)\,dt+(1-f_{1}(0,h\Delta))$
(18) $\cdot\int_{0}^{\infty}F_{1}(k-(i+j);a(t-\Delta))\cdot f_{2}(t-k\Delta;k,h)\,dt\bigg{)}$
@@
4.2. Throughput and Related Trade-offs
An ideal protocol design should select parameters to yield high throughput, fast confirmation, and security against adversary miners. We refer to the average number of blocks mined during the maximum propagation delay, $\lambda\Delta$, as the natural fork number. Throughout the paper, let $\beta$ denote the maximum tolerable fraction of adversarial mining. The natural fork number is bounded by
(19) $\lambda\Delta<\frac{1}{\beta}-\frac{1}{1-\beta}$
due to the fault tolerance (1).

### v5

@@ II-C Adversary and Propagation Delay
A block other than the genesis block is either an A-block or an H-block. The parent of an H-block must be credible when the H-block is mined.
An A-block (resp. H-block) is a mathematical concept that represents a block mined by an adversarial (resp. honest) miner. According to Definitions 4 and 8, an H-block must be higher than all public chains at the time of its mining. There is no such constraint on an A-block. In fact, this is the only constraint on block heights. In Fig. 1, we distinguish each A-block using a pair of horns on top. Blocks 2 and 7 are A-blocks, while all the other blocks are H-blocks.
The propagation delays of blocks and chains depend on the network conditions and may also be manipulated by the adversary. To establish a strong safety guarantee, we adopt the widely used $\Delta$-synchrony model, which is defined by the following constraint:
@@
Let $A_{s,t}$ denote the number of A-blocks mined during $(s,t]$. Let $H_{s,t}$ denote the number of H-blocks mined during $(s,t]$. Evidently, $(A_{0,t}+H_{0,t})_{t\geq 0}$ is a Poisson point process with rate
$\lambda=a+h.$ (2)
@@
As a convention, let $A_{s,t}=H_{s,t}=0$ for all $s>t$. Since the probability that any two or more blocks are mined at exactly the same time is equal to zero, we assume without loss of generality that all blocks are mined at different times. Hence $0=t_{0}<t_{1}<t_{2}<\dots$.
III Model Sufficiency
III-A Sufficiency
@@
The prevailing model in this line of work (see, e.g., [13, 11, 14, 15, 3, 4, 16, 17, 5, 18]) is based on the view-based model. Such a model defines a protocol as a set of algorithms executed in an environment that activates $n$ parties, which may be either honest or corrupted, each with its own view [11, Sec. 2]. The protocol’s execution is no simpler than a system of $n$ interactive Turing machines. The view-based model often assumes discrete time steps in execution.
@@
The model proposed in Sec. II of this paper is entirely self-contained and much simpler than the view-based model. In contrast to previous works as well as [6, 7], our model intentionally excludes miners and their views in definitions. In particular, with well-defined block mining processes, we are not concerned with the number of miners; neither do we keep track of which blocks an individual miner has mined or seen. Instead, our model keeps track of the set of public chains at every point in time, which represents the weakest possible view to any miner. In particular, every honest view is a superset of this weakest view, and conversely, the adversary in the view-based model can ensure at least one honest miner holds that weakest view.
@@
Relative to the view-based model under $\Delta$-synchrony, our model has no loss of generality in the sense that, for a given realization (or sample path) of the A- and H-block mining processes, if the view-based model can produce a certain set of chains, so can our model. This is because both models respect the same constraints for chain creation described in Definitions 8 and 9. For the latency-security analysis, our model is as powerful as the view-based model. To elaborate, consider the event in the view-based model where, first, a chain in an honest view confirms a block on height $h$, and then another chain in an honest view confirms a different block on the same height, creating a safety violation. Our model can produce the same pair of chains, each of which is credible to the weakest view at the time it confirms a conflicting block. Conversely, if our model produces a pair of conflicting credible chains, then those chains can be used to convince two honest parties in the view-based model to confirm different blocks on the same height. Once we realize the role of the weakest view, there is no need to define individual views and their complicated interactions.
III-B Operational Meaning
@@
For concreteness and practicality, we relate the definitions in Sec. II to the operations of an individual miner (or node), who assembles transactions into blocks and then chains by expending proof of work. If the node is honest, it follows the longest-chain fork-choice rule and immediately share new blocks via the peer-to-peer network; otherwise, it can use an arbitrary (possibly randomized) attack strategy. At time $t$, every node has its own view, which includes all transactions, blocks, and chains the node has seen by $t$. The operations of this node is based on the entire history of this node’s view, which can only increase with time.
In general, an honest node will provide goods and services for a transaction only after it commits the transaction. As committing a transaction has no bearing on chain production, nodes are free to choose their individual commit rules. A popular choice is to commit after $k$ block confirmations, where each node can select its own depth number $k$. In operations, after an honest node sees a new block $g$ extending one of the highest chains in the node’s view, the node commits all new transactions in all blocks in chain $g$ as soon as any chain $b$ with height $h_{b}\geq h_{g}+k-1$ becomes a highest chain in its view and has chain $g$ as a prefix. [^2] In case both block $g$ and its child block $c$ enter the node’s view at once, the node can commit transactions in chain $g$ first according to this rule, and then commit additional transactions in chain $c$ in the same manner. We emphasize that block $g$ is $k$-confirmed by chain $b$ by Definition 6. In other words, committed transactions must be $k$-confirmed.
An alternative commit rule is by time of $t$: A node focuses on a block $g$ in its view at (the current) time $s$. If there ever is a time $r\geq t$, such that chain $g$ is a prefix of a highest chain in this node’s view at time $s+r$, then the node commits all new transactions contained in chain $g$. By Definition 5, since this highest chain at time $s+r$ is credible, chain $g$ is confirmed by time of $t$.
With confirmation by either depth or time, once we have a safety guarantee established rigorously for sufficiently confirmed chains in the mathematical model in Sec. II, the honest node enjoys the same operational safety guarantee for every committed transaction. There is in fact the subtlety that block $g$ is anchored to a current view at a point in time chosen by the honest node, so that the adversary cannot directly carry out an adaptive attack of the following form: Mine in private to gain a sufficiently high lead, then pick a block to attack with certainty to cause a violation.
III-C Delay Variations
@@
The seemingly rigid $\Delta$-synchrony model is quite practical. First, “light” nodes do not mine blocks or fork chains, so the delays they perceive are inconsequential to ledger safety as long as honest miners satisfy the delay requirement. Second, if some blocks of honest miners suffer longer delays than $\Delta$, we simply regard those blocks as adversarial blocks to preserve $\Delta$-synchrony. In addition, potential attacks like Erebus [19] and SyncAttack [20] seek to partition the peer-to-peer network. As long as those delay violations are independent of the block mining times, the honest and adversarial mining processes remain to be Poisson. If block propagation delays have a long-tailed distribution [21, 12], we can use a delay bound that applies to a high percentage of blocks. The main results here continue to apply with all out-of-sync mining power treated as being adversarial.
IV Main Results
@@
Recall that $a$ and $h$ represent as the adversarial and honest mining rates, respectively, in the number of blocks mined per second. For confirmation by either time or depth, the following negative result is well understood (see, e.g., [11, 3, 4]).
@@
$\frac{1}{a}<\Delta+\frac{1}{h},$ (3)
then every block is unsafe in the sense that the private-mining attack will violate its safety with probability 1.
@@
IV-A Bounds for Given Confirmation Depth
@@
$b(\alpha,\eta)=\left(\frac{e^{\frac{5}{2}\gamma}\gamma(\eta-\alpha-\alpha\eta)}{\alpha\eta e^{\gamma}-(\alpha+\gamma)(\eta-\gamma)}\right)^{2}+1-\frac{\eta-\alpha-\alpha\eta}{\eta e^{\alpha}},$ (4)
@@
$\begin{split}c(\alpha,\eta)=\left(\ln\left(1+\frac{\gamma}{\alpha}\right)+\ln\left(1-\frac{\gamma}{\eta}\right)-\gamma\right)\cdot\left(1-\ln\left(1+\frac{\gamma}{\alpha}\right)\bigg{/}\ln\frac{\eta e^{\alpha}(e^{\eta}-1)}{\eta+\eta e^{\alpha}(e^{\eta}-1)-\alpha(1+\eta)}\right)^{-1}\end{split}$ (5)
@@
$\gamma(\alpha,\eta)=\frac{2-\alpha+\eta-\sqrt{4+(\alpha+\eta)^{2}}}{2}.$ (6)
The following latency-security trade-off is the first set of non-trivial safety guarantees that apply to all parameter values up to the ultimate fault tolerance.
Theorem 2 (Safety guarantee for given confirmation depth).
@@
$\frac{1}{a}>\Delta+\frac{1}{h}.$ (7)
Given a sufficiently large $s$, the safety of a block extending an $s$-credible chain and being $k$-confirmed can be violated with a probability no higher than
$b(a\Delta,h\Delta)e^{-c(a\Delta,h\Delta)\cdot k}$ (8)
@@
As long as (7) holds, every block is arbitrarily safe after sufficiently long confirmation latency (either in depth or time).
The pair of opposing inequalities (3) and (7) determine the exact boundary between the asymptotically achievable and unachievable latency-security trade-offs. The right hand side of (3) and (7) is exactly the expected inter-arrival time of the following type of blocks:
@@
By definition, every pacer becomes public before the next pacer arrives, and hence all pacers are on different heights. Since the expected pacer inter-arrival time is $\Delta+1/h$, pacers arrive at the rate of $h/(1+h\Delta)$ in the long run. As a consequence, the number of heights H-blocks occupy grows at least at that rate. This approach enables us to use the growth rate of pacers as a lower bound for the height increase of H-blocks.
@@
$f(t)=\frac{1}{h}e^{-h(t-\Delta)}1_{\{t>\Delta\}}.$ (9)
Lemma 4 ([22, Theorem 3.4]).
@@
$\mathcal{E}(r)=\frac{(1-r)(h-a-ha\Delta)}{h-e^{(1-r)a\Delta}(h+a-ar)r}.$ (11)
@@
$e(i)$ $=\frac{1}{i!}\mathcal{E}^{(i)}(0),\qquad i=0,1,\dots$ (12)
@@
$\delta$ $=\left(1-a\frac{1+h\Delta}{h}\right)^{+}e^{-a\Delta}.$ (14)
@@
$F^{(k)}(n)=1-(1-\delta)\left(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\right)^{n+1}\bar{F}_{b}(n;k,(\epsilon+\delta-\epsilon\delta))$ (15)
@@
$F(n)$ $=\lim_{k\to\infty}F^{(k)}(n)$ (16)
$=1-(1-\delta)\left(\frac{\epsilon}{\epsilon+\delta-\epsilon\delta}\right)^{n+1}.$ (17)
@@
Given a sufficiently large $s$, the safety of a block that extends an $s$-credible chain and is confirmed by the depth of $k$ can be violated with a probability no higher than
$\inf_{n}\bigg{\{}2-$ $\sum_{i=0}^{k-n}e(i)\sum_{j=0}^{k-n-i}e(j)\int_{0}^{\infty}F_{1}(k-n-i-j;a(t+3\Delta))\cdot f_{2}(t-(k)\Delta;k,h)\,dt-F^{(k)}(n-1)\bigg{\}}$ (18)
@@
Definition 12 (Private-mining attack on a target block).
The adversary delays the propagation of every H-block maximally (for $\Delta$ seconds). Starting from time 0, the adversary always mines in private to extend the highest chain which does not include the target block.
Recall that Definition 7 allows nonpublic credible chains to cause a violation. However, operationally, the adversary can carry out a private-mining attack and make the private chain public as soon as both of the following conditions hold: 1. the block has been confirmed by a chain of H-blocks; and 2. the private chain which does not contain the block is credible. In other words, as long as a safety violation occurs by Definition 7, the adversary creates a violation operationally.
@@
Given a sufficiently large $s$, the private-mining attack violates the safety of a target block which extends an $s$-credible chain with at least the following probability:
@@
IV-B Throughput and Related Trade-offs
An ideal protocol design should select parameters to yield high throughput, fast confirmation, and security against adversarial miners. We refer to the average number of blocks mined during the maximum propagation delay, $\lambda\Delta$, as the natural fork number. Throughout the paper, let $\beta$ denote the maximum tolerable fraction of adversarial mining. To operate within fault tolerance (1), the natural fork number must be bounded by
$\lambda\Delta<\frac{1}{\beta}-\frac{1}{1-\beta}.$ (20)
We remark that it is more convenient to discuss throughput in terms of $\beta$ and the total mining rate $\lambda$ in lieu of the adversarial and honest mining rates ($a$ and $h$).